# Capturing Turbulence with Numerical Dissipation: a Simple Dynamical Model for Unresolved Turbulence in Hydrodynamic Simulations

### v1

@@ 1 Introduction
To this aim, we propose a hybrid method that combines features of both explicit and implicit LES, which we dub as semi-implicit LES (SLES). In this method, the small-scale turbulent energy is followed explicitly as in standard LES. However, the turbulence source term is computed implicitly by capturing locally dissipated resolved kinetic energy and converting it into subgrid turbulence. This is motivated by the above feature of turbulent flows that the local dissipation rate is set by the energy cascade rate, which constitutes the basis of ILES. The model is straightforward to implement as it relies on the existing infrastructure of ILES codes, while also making the local estimate of unresolved turbulence readily available as in explicit LES.
@@ 3 Decaying supersonic turbulence
In this section, we use a decaying supersonic isothermal turbulence test to calibrate and investigate the performance of our new Semi-implicit Large Eddy Simulation (SLES) model. To this end, we use a high-resolution simulation to predict the distribution of turbulent energy on small scales. This distribution is then compared with low-resolution re-simulations of the same initial conditions where the small-scale turbulence is predicted using the SLES model. As we will show, despite its simplicity, the SLES model is able to capture multiple non-trivial features of the DNS simulations, such as the time evolution of the mean turbulent energy, its dependence on scale, and the correlation between gas density and turbulent velocity.
@@ 3.1 Numerical setup and analysis
Figure 2: Gas density (gray) and small-scale turbulent energy slices (orange) in the coarse-grained high-resolution simulation (DNS; the top set of panels) and low-resolution re-simulations of the same ICs with the SLES model (the bottom set of panels). The maps are shown at $t=2/3\;t_{\rm cross,0}$. The leftmost panel in the top row shows the density slice at the native resolution of the DNS, while the other columns show coarse-grained values, with density being volume-weighted and $e_{\rm turb}$ computed from Equation (8); the corresponding grid size is indicated at the top. The bottom panels show slices at the same timestep from low-resolution re-simulations using corresponding grid sizes where $e_{\rm turb}$ is predicted using the SLES model. Thus, in total, four simulations are shown in the plot as indicated by the green shaded background: one DNS and three SLES re-simulations. Low-resolution SLES re-simulations produce smoother solutions than the DNS but they capture main global features of the small-scale turbulence, in particular, the correlation with gas density.
@@
The results of such low-resolution re-simulations with the SLES model are shown in the bottom set of panels. These low-resolution re-simulations reproduce the overall features of the direct simulation such as the locations of the high-density structures and the correspondence between high-density and high-$e_{\rm turb}$ regions. The solution is significantly smoother than in the DNS owing to higher numerical viscosity, i.e., larger numerical dissipation scales. The SLES model also captures the increase of the mean $e_{\rm turb}$ at a lower resolution.
@@ 3.2 Model calibration
Figure 3: Comparison of the global mean small-scale turbulent energy, $E_{\rm turb}\equiv\langle e_{\rm turb}\rangle_{\rm box}$, in coarse-grained DNS (dashed line) and low-resolution SLES re-simulations with different values of the dissipation parameter, $C_{\epsilon}$ (colored lines). For this comparison, we used the grid size of 323, i.e., the cell size $\Delta\equiv L_{\rm box}/32$. In the DNS, $e_{\rm turb}$ is computed following Equation (8) while in the low-resolution resimulations it is tracked by the SLES model. The mean $E_{\rm turb}$ is normalized to the total initial kinetic energy, i.e., $E_{\rm kin,0}\equiv E_{\rm turb}(<L_{\rm box})$ at $t=0$. The time-axis is normalized by the initial box-crossing time of turbulence, $t_{\rm cross,0}\equiv 0.5\;L_{\rm box}/\sigma_{0}$, or $\approx 0.058$ sound-crossing times for the initial Mach number of $\mathcal{M}\approx 8.6$. The value of $C_{\epsilon}=1.6$ provides the best match to the global $E_{\rm turb}$ and therefore we will use it as our fiducial choice in the rest of the paper. The agreement with the DNS can be further improved by making $C_{\epsilon}$ dependent on the local subgrid Mach number as shown with the thin line and described in the text.
@@ 3.3.1 Mean small-scale turbulent energy
Figure 4: Comparison of coarse-grained DNS (dashed lines) with the SLES re-simulations at corresponding resolutions (colored lines). The top panel shows the evolution of the global Mach number, $\mathcal{M}=\sqrt{2E_{\rm kin}}$ in the adopted units ($\langle\rho\rangle\equiv 1$). The bottom panel shows the mean small-scale turbulent energy, $E_{\rm turb}\equiv\langle e_{\rm turb}\rangle_{\rm box}$, normalized by the initial total kinetic energy in the box, $E_{\rm kin,0}$. The SLES model captures the As shown in Appendix A, $E_{\rm turb}$ is underestimated for 1283 due to lack of convergence on these small scales; however, we still show the results for this grid to demonstrate that the scaling of $E_{\rm turb}$ with the grid size remains the same in SLES.
@@
Another test of the model is its performance on different spatial scales. In the previous section, the model was calibrated on a specific scale, $L_{\rm box}/32$, corresponding to a $32^{3}$ grid. The bottom panel of Figure 4 shows the model’s performance on different scales, corresponding to $16^{3}$, $32^{3}$, $64^{3}$, and $128^{3}$ grids. The dashed lines show the DNS coarse-grained on these scales, while the solid lines are low-resolution re-simulations with the SLES model using corresponding grids.
@@
One noticeable deviation of the SLES re-simulations from the DNS is the slightly higher global Mach number and mean $E_{\rm turb}$ at $t\sim 1\text{--}3\times t_{\rm cross,0}$. As was shown in the previous section, such a feature can be suppressed by modifying the model for the subgrid turbulence decay (see Figure 3). However, this feature might also be due to the initial transient associated with the smoothing of initial sharp cell-to-cell density and velocity gradients resulting from the coarse-graining high-resolution ICs. We checked that this feature does not depend on how the subgrid $e_{\rm turb}$ is initialized. In these tests we initialize as $e_{\rm turb}=0$ throughout the box; however, if we start from the $e_{\rm turb}$ equal to the coarse-grained ICs values using Equation (8), this feature still persists. Such a transient can be avoided by testing the model in a continuously driven turbulence box setup, which we leave to a separate study.
@@ 3.3.2 Turbulent velocity–density correlation
Figure 6: Top panel: Comparison of the $\rho$$\sigma$ relation between the coarse-grained DNS (dashed black lines) and low-resolution SLES re-simulation (solid blue lines). Bottom set of panels: dependence of scatter in $\sigma$ on gas compression rate (negative divergence of the resolved velocity field, $-\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}$) at three fixed densities indicated with green lines in the top panel and labeled in the corner of each panel. Each of the panels shows cells with densities bound by 1.5 times higher and 1.5 times lower than these values. For reference, the dotted red lines show the medians from the low-resolution resimulations with an explicit LES model proposed by schmidt11. The SLES model qualitatively reproduces the trend of $\sigma$ with $\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}$ although it does show a steeper increase at high compression rates. This behavior, however, is analogous to that in a more complex explicit LES model.
As Figure 2 showed, local turbulent energy on small scales is correlated with the gas density. Figure 5 demonstrated this correlation quantitatively for both the DNS coarse-grained using a $32^{3}$ grid (left column) and a SLES resimulation at the corresponding resolution (right column). The rows correspond to different times, from top to bottom: 1, 3, and $5$ global turbulence crossing time of the simulation box, $t_{\rm cross,0}$. The range of the $y$-axes changes from top to bottom to compensate for the decrease in the normalization of this relation due to turbulence decay; to help visualize this decreasing trend, the red line is kept the same in all panels (see below).
@@ 3.3.3 Power spectra
Figure 7: Power spectra in DNS (dashed lines) and SLES re-simulations (solid colored lines): velocity (${\boldsymbol{u}}$; top), density ($\rho$; middle), and small-scale turbulent energy ($e_{\rm turb}$; bottom). The velocity power spectrum is compensated by $k^{-5/3}$, while $\rho$ and $e_{\rm turb}$ spectra are compensated by $k^{-1.1}$. Gray dashed lines show the results from DNS at native 10243 resolution, while thin colored dashed lines show coarse-grained statistics on the grids corresponding to the SLES re-simualtions. The wave number, $k$, is normalized by the minimal value corresponding to the box size, $k_{\rm min}=2\pi/L_{\rm box}$. Vertical dotted lines show Nyquist frequencies for low-resolution grids, i.e., $k_{\rm Ny}=\pi N_{\rm grid}/L_{\rm box}$.
@@
As the spatial resolution is degraded in SLES re-simulations, the kinetic energy is dissipated on larger scales (see the colored lines in the top panel of Figure 7). For reference, the vertical dotted lines in the figure show the Nyquist wave numbers corresponding to 2 grid cells at a given resolution: $k_{\rm Ny}=\pi N_{\rm grid}/L_{\rm box}$. The offset of numerical dissipation to larger and larger scales leads to the distribution of density becoming progressively smoother (see the bottom set of maps in Figure 2).
@@ 4.2 Distribution of small-scale turbulence
One of the physical reasons why these distributions are similar in dense regions is that they are largely shaped by compressive heating (via the $PdV$ work) and decay of turbulence (semenov16; semenov17), and these processes that are modeled in exactly the same way in both models (see Section 2). The evolution of gas in this part of the $n$$\sigma_{\rm tot}$ plane is, in fact, analogous to the evolution of gas in our supersonic developed turbulence tests in Figure 5: at high $\sigma_{\rm tot}$, gas density on average increases, implying $\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}<0$, while at low $\sigma_{\rm tot}$, gas density decreases and $\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}>0$, producing clock-wise circulation of gas in that plane (see Figures 4 and 6 in semenov17). Thus, as long as subgrid turbulence is produced at a reasonable level in diffuse gas in diffuse gas, gas compression during the formation of dense regions will bring it to the values set by the interplay between the $PdV$ work and turbulence decay.
In diffuse gas at the disk outskirts, in contrast, the values of $\sigma_{\rm turb}$ are different (recall Figure 8). However, in these regions, turbulence is strongly subsonic, and $\sigma_{\rm tot}$ is dominated by the thermal sound speed. As a result, in these regions, small-scale turbulence does not affect gas dynamics, as its contribution to the total pressure is small.
@@ 4.3 Global and kiloparsec-scale star formation
Figure 10: Evolution of global star formation properties in our galaxy simulations with two different models for unresolved turbulence. The panels show, from top to bottom: total SFR, total mass of star-forming gas, its average density, and locall star formation efficiency per freefall time, $\epsilon_{\rm ff}$. Solid lines show the tracks smoothed on a 100 Myr timescale using a Gaussian filter. The SLES model results in $\sim 2$ times lower SFR than the explicit LES simulation. This difference is mainly due to the difference in the predicted local $\epsilon_{\rm ff}$, while the total mass and average density of star-forming gas are similar between the runs.
@@ 5 Discussion
One example is the disk outskirts and very center, where SLES model predicts higher turbulent velocities than the explicit LES model (Section 4). One reason for this discrepancy is the difference in the treatment of shearing flows. Our explicit model adopts a shear-improved LES closure where the bulk flow is filtered out and only the fluctuating part of the velocity field is interpreted as a part of the turbulent cascade (leveque07, see also Section 2.2). To avoid spurious production of $e_{\rm turb}$ in stable shearing flows, such as differentially rotating disks, one needs to adopt a similar treatment in the SLES model. For example, instead of putting all dissipated energy into $e_{\rm turb}$ one can make it a function of the difference between the local filtered and total flows, such that in purely shearing flows kinetic energy is dissipated directly into heat. We leave the investigation of such models to future work.

### v2

@@ 1 Introduction
To this aim, we propose a hybrid method that combines features of both explicit and implicit LES, which we dub as semi-implicit LES (SLES). In this method, the small-scale turbulent energy is followed explicitly, as is the case with standard LES. However, the turbulence source term is computed implicitly by capturing locally dissipated resolved kinetic energy and converting it into subgrid turbulence. This is motivated by the above feature of turbulent flows that the local dissipation rate is set by the energy cascade rate, which constitutes the basis of ILES. The model is straightforward to implement as it relies on the existing infrastructure of ILES codes, while also making the local estimate of unresolved turbulence readily available as in explicit LES.
@@ 3 Decaying supersonic turbulence
In this section, we use a decaying supersonic isothermal turbulence test to calibrate and investigate the performance of our new Semi-implicit Large Eddy Simulation (SLES) model. To this end, we use a high-resolution simulation to predict the distribution of turbulent energy on small scales. This distribution is then compared with low-resolution resimulations of the same initial conditions where the small-scale turbulence is predicted using the SLES model. As we will show, despite its simplicity, the SLES model is able to capture multiple non-trivial features of the DNS simulations, such as the time evolution of the mean turbulent energy, its dependence on scale, and the correlation between gas density and turbulent velocity.
@@ 3.1 Numerical setup and analysis
Figure 2: Gas density (gray) and small-scale turbulent energy slices (orange) in the coarse-grained high-resolution simulation (DNS; the top set of panels) and low-resolution resimulations of the same ICs with the SLES model (the bottom set of panels). The maps are shown at $t=2/3\;t_{\rm cross,0}$. The leftmost panel in the top row shows the density slice at the native resolution of the DNS, while the other columns show coarse-grained values, with density being volume-weighted and $e_{\rm turb}$ computed from Equation (8); the corresponding grid size is indicated at the top. The bottom panels show slices at the same timestep from low-resolution resimulations using corresponding grid sizes where $e_{\rm turb}$ is predicted using the SLES model. Thus, in total, four simulations are shown in the plot as indicated by the green shaded background: one DNS and three SLES resimulations. Low-resolution SLES resimulations produce smoother solutions than the DNS but they capture main global features of the small-scale turbulence, in particular, the correlation with gas density.
@@
The results of such low-resolution resimulations with the SLES model are shown in the bottom set of panels. These low-resolution resimulations reproduce the overall features of the direct simulation such as the locations of the high-density structures and the correspondence between high-density and high-$e_{\rm turb}$ regions. The solution is significantly smoother than in the DNS owing to higher numerical viscosity, i.e., larger numerical dissipation scales. The SLES model also captures the increase of the mean $e_{\rm turb}$ at a lower resolution.
@@ 3.2 Model calibration
Figure 3: Comparison of the global mean small-scale turbulent energy, $E_{\rm turb}\equiv\langle e_{\rm turb}\rangle_{\rm box}$, in coarse-grained DNS (dashed line) and low-resolution SLES resimulations with different values of the dissipation parameter, $C_{\epsilon}$ (colored lines). For this comparison, we used the grid size of 323, i.e., the cell size $\Delta\equiv L_{\rm box}/32$. In the DNS, $e_{\rm turb}$ is computed following Equation (8) while in the low-resolution resimulations, it is tracked by the SLES model. The mean $E_{\rm turb}$ is normalized to the total initial kinetic energy, i.e., $E_{\rm kin,0}\equiv E_{\rm turb}(<L_{\rm box})$ at $t=0$. The time-axis is normalized by the initial box-crossing time of turbulence, $t_{\rm cross,0}\equiv 0.5\;L_{\rm box}/\sigma_{0}$, or $\approx 0.058$ sound-crossing times for the initial Mach number of $\mathcal{M}\approx 8.6$. The value of $C_{\epsilon}=1.6$ provides the best match to the global $E_{\rm turb}$ and therefore we will use it as our fiducial choice in the rest of the paper. The agreement with the DNS can be further improved by making $C_{\epsilon}$ dependent on the local subgrid Mach number as shown with the thin line and described in the text.
@@ 3.3.1 Mean small-scale turbulent energy
Figure 4: Comparison of coarse-grained DNS (dashed lines) with the SLES resimulations at corresponding resolutions (colored lines). The top panel shows the evolution of the global Mach number, $\mathcal{M}=\sqrt{2E_{\rm kin}}$ in the adopted units ($\langle\rho\rangle\equiv 1$). The bottom panel shows the mean small-scale turbulent energy, $E_{\rm turb}\equiv\langle e_{\rm turb}\rangle_{\rm box}$, normalized by the initial total kinetic energy in the box, $E_{\rm kin,0}$. The SLES model captures both the evolution of $E_{\rm turb}$ and its scale dependence. As shown in Appendix A, $E_{\rm turb}$ is underestimated for 1283 due to lack of convergence on these small scales; however, we still show the results for this grid to demonstrate that the scaling of $E_{\rm turb}$ with the grid size remains the same in SLES.
@@
Another test of the model is its performance on different spatial scales. In the previous section, the model was calibrated on a specific scale, $L_{\rm box}/32$, corresponding to a $32^{3}$ grid. The bottom panel of Figure 4 shows the model’s performance on different scales, corresponding to $16^{3}$, $32^{3}$, $64^{3}$, and $128^{3}$ grids. The dashed lines show the DNS coarse-grained on these scales, while the solid lines are low-resolution resimulations with the SLES model using corresponding grids.
@@
One noticeable deviation of the SLES resimulations from the DNS is the slightly higher global Mach number and mean $E_{\rm turb}$ at $t\sim 1\text{--}3\times t_{\rm cross,0}$. As was shown in the previous section, such a feature can be suppressed by modifying the model for the subgrid turbulence decay (see Figure 3). However, this feature might also be due to the initial transient associated with the smoothing of initial sharp cell-to-cell density and velocity gradients resulting from the coarse-graining high-resolution ICs. We checked that this feature does not depend on how the subgrid $e_{\rm turb}$ is initialized. In these tests we initialize as $e_{\rm turb}=0$ throughout the box; however, if we start from the $e_{\rm turb}$ equal to the coarse-grained ICs values using Equation (8), this feature still persists. Such a transient can be avoided by testing the model in a continuously driven turbulence box setup, which we leave to a separate study.
@@ 3.3.2 Turbulent velocity–density correlation
Figure 6: Top panel: Comparison of the $\rho$$\sigma$ relation between the coarse-grained DNS (dashed black lines) and low-resolution SLES resimulation (solid blue lines). Bottom set of panels: dependence of scatter in $\sigma$ on gas compression rate (negative divergence of the resolved velocity field, $-\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}$) at three fixed densities indicated with green lines in the top panel and labeled in the corner of each panel. Each of the panels shows cells with densities bound by 1.5 times higher and 1.5 times lower than these values. For reference, the dotted red lines show the medians from the low-resolution resimulations with an explicit LES model proposed by schmidt11. The SLES model qualitatively reproduces the trend of $\sigma$ with $\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}$ although it does show a steeper increase at high compression rates. This behavior, however, is analogous to that in a more complex explicit LES model.
As Figure 2 showed, local turbulent energy on small scales is correlated with the gas density. Figure 5 demonstrated this correlation quantitatively for both the DNS coarse-grained using a $32^{3}$ grid (left column) and an SLES resimulation at the corresponding resolution (right column). The rows correspond to different times, from top to bottom: 1, 3, and $5$ global turbulence crossing time of the simulation box, $t_{\rm cross,0}$. The range of the $y$-axes changes from top to bottom to compensate for the decrease in the normalization of this relation due to turbulence decay; to help visualize this decreasing trend, the red line is kept the same in all panels (see below).
@@ 3.3.3 Power spectra
Figure 7: Power spectra in DNS (dashed lines) and SLES resimulations (solid colored lines): velocity (${\boldsymbol{u}}$; top), density ($\rho$; middle), and small-scale turbulent energy ($e_{\rm turb}$; bottom). The velocity power spectrum is compensated by $k^{-5/3}$, while $\rho$ and $e_{\rm turb}$ spectra are compensated by $k^{-1.1}$. Gray dashed lines show the results from DNS at native 10243 resolution, while thin colored dashed lines show coarse-grained statistics on the grids corresponding to the SLES resimulations. The wave number, $k$, is normalized by the minimal value corresponding to the box size, $k_{\rm min}=2\pi/L_{\rm box}$. Vertical dotted lines show Nyquist frequencies for low-resolution grids, i.e., $k_{\rm Ny}=\pi N_{\rm grid}/L_{\rm box}$.
@@
As the spatial resolution is degraded in SLES resimulations, the kinetic energy is dissipated on larger scales (see the colored lines in the top panel of Figure 7). For reference, the vertical dotted lines in the figure show the Nyquist wave numbers corresponding to 2 grid cells at a given resolution: $k_{\rm Ny}=\pi N_{\rm grid}/L_{\rm box}$. The offset of numerical dissipation to larger and larger scales leads to the distribution of density becoming progressively smoother (see the bottom set of maps in Figure 2).
@@ 4.2 Distribution of small-scale turbulence
One of the physical reasons why these distributions are similar in dense regions is that they are largely shaped by compressive heating (via the $PdV$ work) and decay of turbulence (semenov16; semenov17), and these processes that are modeled in exactly the same way in both models (see Section 2). The evolution of gas in this part of the $n$$\sigma_{\rm tot}$ plane is, in fact, analogous to the evolution of gas in our supersonic developed turbulence tests in Figure 5: at high $\sigma_{\rm tot}$, gas density on average increases, implying $\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}<0$, while at low $\sigma_{\rm tot}$, gas density decreases and $\boldsymbol{\nabla}\boldsymbol{\cdot}{\boldsymbol{u}}>0$, producing clock-wise circulation of gas in that plane (see Figures 4 and 6 in semenov17). Thus, as long as subgrid turbulence is produced at a reasonable level in diffuse gas, gas compression during the formation of dense regions will bring it to the values set by the interplay between the $PdV$ work and turbulence decay.
In diffuse gas at the disk outskirts, in contrast, the values of $\sigma_{\rm turb}$ are different (recall Figure 8). However, in these regions, turbulence is strongly subsonic, and $\sigma_{\rm tot}$ is dominated by the thermal sound speed, which creates the lower envelope of the $n$$\sigma_{\rm tot}$ distribution at $n<10\;{\rm cm^{-3}}$. As a result, at these densities, small-scale turbulence does not affect gas dynamics, as its contribution to the total pressure is small.
@@ 4.3 Global and kiloparsec-scale star formation
Figure 10: Evolution of global star formation properties in our galaxy simulations with two different models for unresolved turbulence. The panels show, from top to bottom: the total SFR, the total mass of star-forming gas, its average density, and local star formation efficiency per freefall time, $\epsilon_{\rm ff}$. Solid lines show the tracks smoothed on a 100 Myr timescale using a Gaussian filter. The SLES model results in $\sim 2$ times lower SFR than the explicit LES simulation. This difference is mainly due to the difference in the predicted local $\epsilon_{\rm ff}$, while the total mass and average density of star-forming gas are similar between the runs.
@@ 5 Discussion
One example is the disk outskirts and very center, where the SLES model predicts higher turbulent velocities than the explicit LES model (Section 4). One reason for this discrepancy is the difference in the treatment of shearing flows. Our explicit model adopts a shear-improved LES closure where the bulk flow is filtered out and only the fluctuating part of the velocity field is interpreted as a part of the turbulent cascade (leveque07, see also Section 2.2). To avoid spurious production of $e_{\rm turb}$ in stable shearing flows, such as differentially rotating disks, one needs to adopt a similar treatment in the SLES model. For example, instead of putting all dissipated energy into $e_{\rm turb}$ one can make it a function of the difference between the local filtered and total flows, such that in purely shearing flows kinetic energy is dissipated directly into heat. We leave the investigation of such models to future work.